# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v3

@@ 4.1 Experiment Setup
Datasets and Baselines. For pre-training CAPE, PatchTST, and PEM, we manually collected 17 distinct weekly-sampled diseases from Project Tycho (van Panhuis et al., 2018). For evaluation, we utilize five downstream datasets covering various diseases and locations: ILI USA (Centers for Disease Control and Prevention, 2023a), ILI Japan (National Institute of Infectious Diseases, 2023), COVID-19 USA (Dong et al., 2020), Measles England (Lau et al., 2020a), and Dengue across countries (OpenDengue, 2023). Additionally, RSV (Centers for Disease Control and Prevention, 2023c) and MPox (Centers for Disease Control and Prevention, 2023b) infections in the US are used to test zero-shot performance. More details can be found in Appendix A.4. We adopt baseline models from the comprehensive EpiLearn toolkit (Liu et al., 2024a), comparing CAPE against two categories: non-pretrained and pre-trained models. Non-pretrained baselines include classical statistical models such as ARIMA (Panagopoulos et al., 2021), configured in an online forecasting setting (Pham et al., 2022; Abdulmajeed et al., 2020) where parameters are updated with each new sample. We also consider deep learning models, including RNN-based approaches like LSTM and GRU (Wang et al., 2020; Natarajan et al., 2023), the linear model DLinear (Zeng et al., 2023), and transformer-based architectures (Wu et al., 2021; Zhou et al., 2021, 2022). Pre-trained baselines include state-of-the-art models such as PatchTST (Nie et al., 2022), PEM (Kamarthi & Prakash, 2023), and the time series foundation model MOMENT (Goswami et al., 2024). Further experimental details are provided in Appendix A.5.
@@ 4.2 Fine-Tuning (Full-Shot Setting)
For non-pre-trained models, we train the entire model on the training split, while for pre-trained models, we fine-tune on downstream datasets by transferring the task-specific head $h_{\psi}$ from pre-training to the forecasting task. We evaluate short-term and long-term performance by reporting MSE across horizons from 1 to 16. From Table 1, we observe: (a) CAPE achieves the best average MSE across all downstream datasets. It outperforms the best baseline by 9.91% on average and up to 14.85%. On the COVID dataset, CAPE performs best across all horizons, showing effectiveness on novel diseases. (b) Models like PEM, PatchTST, and MOMENT consistently rank second or third on 4 out of 5 downstream datasets. The best pre-trained model (excluding CAPE) outperforms the best non-pre-trained model by 6.223% on average. Among them, PatchTST has the highest average performance, surpassing PEM by 5.51% and MOMENT by 10.45%. Additionally, PEM outperforms MOMENT by 4.86%, indicating the importance of epidemic-specific pre-training. (c) Informer consistently outperforms Autoformer and Fedformer by 24.40% and 17.90% respectively, due to its sparse attention mechanism that reduces overfitting. Informer also surpasses Dlinear by 1.90%, suggesting that careful selection of model size and parameters is crucial for optimal performance. (d) Furthermore, environment modeling proves valuable, as CAPE consistently outperforms PatchTST, which shares a similar design. While both models are pre-trained on the epidemic-specific datasets, CAPE surpasses PatchTST by 11.13%.
4.3 Few-Shot Performance, Zero-Shot Performance, and Ablation Study
Few-Shot Forecasting. In real-world scenarios, predicting outbreaks of diseases unknown or in new locations is challenging for purely data-driven models due to limited initial data. Thus, few-shot or zero-shot forecasting capabilities are essential for epidemic models. To simulate a few-shot scenario, we reduce the original training data from 100% to [20%, 40%, 60%, 80%]. We report the average MSE across 1 to 16 time steps. From Table 11, we make the following observations: (a) With an increasing volume of training materials, the performance of all models consistently improves. (b) CAPE achieves the best performance in most scenarios, demonstrating the superior few-shot ability. (c) Compared with models pre-trained on epidemic-specific datasets, Dlinear failed to achieve better performance when only 20% of training data is available. However, Dlinear is able to outperform MOMENT on ILI USA and Measles datasets when both models are trained or fine-tuned using 20% training data, which indicates the importance of pre-training. (d) Though CAPE achieves the best average performance on the ILI USA dataset when the training material is reduced, it achieves a good performance in short-term forecasting from 1 to 4 weeks (see Appendix A.9).
@@ 4.3 Few-Shot Performance, Zero-Shot Performance, and Ablation Study
Dataset $\Delta$ (%) CAPE PatchTST PEM MOMENT
ILI USA 9.26% 0.147 0.164 0.162 0.549
ILI Japan 17.06% 0.705 0.907 0.850 2.062
Measles 3.97% 0.145 0.167 0.159 0.533
MPox 20.00% 0.0004 0.0005 0.0005 0.0013
Dengue (mixed) 10.17% 0.371 0.427 0.413 1.624
RSV 26.06% 0.834 1.128 1.260 1.849
Covid (daily interval) 13.80% 5.173 6.001 6.320 18.881
Table 4: Ablation study.
Dataset Model H=1 H=2 H=4 H=8 H=16 Avg
ILI USA CAPE 0.155 0.200 0.270 0.404 0.516 0.309
w/o Env 0.326 0.448 0.508 0.642 0.735 0.532
w/o Contrast 0.174 0.241 0.335 0.492 0.570 0.363
w/o Pretrain 0.158 0.202 0.283 0.408 0.545 0.319
Measles CAPE 0.069 0.096 0.155 0.280 0.743 0.269
w/o Env 0.083 0.111 0.168 0.407 0.755 0.304
w/o Contrast 0.090 0.124 0.276 0.431 0.801 0.344
w/o Pretrain 0.074 0.113 0.223 0.402 0.816 0.326
Dengue CAPE 0.218 0.301 0.540 1.193 2.210 0.892
w/o Env 0.232 0.316 0.484 1.089 3.622 1.149
w/o Contrast 0.198 0.273 0.460 1.128 3.329 1.078
w/o Pretrain 0.210 0.276 0.449 1.115 3.759 1.162
Zero-Shot Forecasting. To further demonstrate the potential of our model, we evaluate CAPE in a zero-shot setting. Specifically, for transformer-based models, we retain the pre-training head and freeze all parameters during testing. All models are provided with a short input sequence of 12 time steps and tasked with predicting infections for the next 4 time steps. From Table 4, we make the following observations: (a) CAPE outperforms baselines across all downstream datasets, showing superior zero-shot forecasting ability. (b) Models pre-trained on epidemic-specific datasets achieve better performance compared to those pre-trained without epidemic-specific data (MOMENT). This indicates the necessity of choosing domain-specific materials for pre-training.
Ablation Study. We conducted an ablation study to assess CAPE’s components (Table 4). Replacing environment estimators with non-disentangling self-attention layers consistently worsened performance across all datasets, notably increasing ILI USA’s MSE from 0.309 to 0.532. Similarly, removing contrastive loss while retaining environment estimators raised Measles’ MSE from 0.269 to 0.344, with smaller increases for ILI USA and Dengue. Training CAPE directly on downstream datasets without pre-training also decreased performance, with MSE rising to 0.319 (ILI USA), 0.326 (Measles), and 1.162 (Dengue), though less than removing environment estimation. These results indicate that all CAPE components are essential for optimal forecasting.
4.4 Transferability
Cross-Location: We include measles data from the USA in the pre-training dataset. To evaluate our model’s ability to adapt to cross-region data, we incorporate measles outbreak data from the UK into the downstream datasets. As shown in Table 4, the pre-trained CAPE outperforms the non-pre-trained version by 17.48%. While we pre-train our model with influenza data from the USA, the zero-shot evaluation on the influenza outbreak in Japan also shows superior performance, underscoring the crucial role of pre-training in enabling generalization to novel regions. Cross-Disease: While we include various types of diseases in our pre-training dataset, novel diseases including Dengue (non-respiratory) and COVID-19 that are unseen in the pre-training stage are incorporated during the downstream evaluation. The ability of our model to adapt to novel diseases is proven compared to the version not pre-trained on the Dengue dataset (Table 4), improving which by 23.24%, as well as the superior zero-shot performance on the COVID dataset (Table 4), which surpasses the MOMENT that is not pre-trained on other diseases by 72.60%. Cross-Interval: While we only pre-train using weekly-sampled data, our model outperformed the non-pre-trained version on the irregularly sampled Dengue dataset and on the daily-sampled COVID-19 dataset, demonstrating robustness to different time intervals.
@@ 4.5 Deeper Analysis
Impact of Pre-Training Epochs: Evaluating four downstream datasets (Figure 5), we find that increasing pre-training epochs consistently improves performance on Measles and COVID datasets but degrades it for ILI USA. Additionally, models with more environment states $K$ perform better as pre-training epochs increase. Impact of Pre-Training Materials: We examine potential biases in our pre-training dataset by splitting it into respiratory and non-respiratory diseases. As shown in Figure 4, with similar volumes of pre-train data, the model performs better when the tested disease types align with the pre-train data. However, the size of the pre-train material has a stronger impact. Impact of Pre-Training Material Scale: To explore how the pre-training material scale affects downstream performance, we scaled the original pre-training dataset and test on downstream datasets. As shown in Figure 6, a sudden performance boost is observed at around a 60% reduction for both Measles and Dengue datasets.
Tackling Distribution Shift: We define distribution shifts as changes in infection patterns from training to test data. We compute the Central Moment Discrepancy (CMD)(Zellinger et al., 2017) between training and test distributions for each disease. As shown in Figure 7, CAPE with environment estimation achieves the lowest CMD scores, highlighting its effectiveness in mitigating distribution shifts. Disentangling Disease Dynamics: We assess CAPE’s ability to capture intrinsic disease dynamics by extracting latent embeddings and computing the Davies-Bouldin Index (DBI) across different dataset pairs. As shown in A.13, CAPE consistently yields lower DBI scores than PatchTST, indicating better separation of disease-specific patterns from environmental influences.
Figure 5: Downstream performance vs. environments distribution. More results in A.10.
Figure 6: Downstream performance across pre-training ratios. More datasets are in  A.11.
Figure 7: CMD scores w/ and w/o environment estimate between train/test sets. (A.12)
@@ 5 Conclusion
We present CAPE, a covariate-adjusted pre-training framework for epidemic forecasting that captures environment-invariant disease dynamics by combining pre-training with implicit environment modeling grounded in causal analysis. By learning latent environments and applying backdoor adjustment, CAPE reduces distribution shifts. Extensive experiments across diseases, regions, and scenarios demonstrate strong generalization, with consistent gains in few-shot and zero-shot settings. Future directions can include incorporating uncertainty measurement, scaling to larger datasets, and providing interpretable environment representations.
@@ A.1 Potential Impact
Early Insights. We provide novel insights into how pre-training and environment modeling improve epidemic forecasting. Our results demonstrate that pre-training significantly enhances model accuracy, with gains increasing as more pre-training data is incorporated. This finding paves the way for future research to develop foundational models in epidemic analysis using larger datasets. Additionally, we confirm the importance of accounting for both inherent disease dynamics and environmental factors to achieve robust forecasting performance.
@@ A.2 Limitations
While CAPE shows strong performance, it has several limitations. First, it does not currently incorporate uncertainty estimation, which is important for risk-aware decision-making. Second, its effectiveness may be constrained by the scale and diversity of available data. Lastly, the learned environment representations lack interpretability, limiting their transparency and potential for insight in public health contexts.
Though not explicitly designed for uncertainty, CAPE supports Dropout Injection (Ledda et al., 2023) at inference. We use 25 stochastic forward passes on the ILI USA dataset to compute 95% confidence intervals:
Table 5: CAPE Forecasting Results with 95% Confidence Intervals
Horizon CAPE 95% CI
1 $0.174\pm 0.0036$
2 $0.235\pm 0.0016$
4 $0.293\pm 0.0012$
8 $0.433\pm 0.0010$
A.3 Theoretical Analysis
A.3.1 Derivation for do-operation
We derive a tractable form for $\log p_{\Theta}(\hat{Y}|do(X))$, leveraging two rules of do-calculus.
Do-Calculus Rules.
Consider a causal DAG $\mathcal{A}$ with nodes $B$, $D$, and $Z$. Let $\mathcal{A}_{B}$ denote the intervened graph by removing all arrows entering $B$, and $\mathcal{A}^{B}$ the graph by removing all arrows leaving $B$. The rules are:
1. Action/Observation Exchange:
$P(D|do(B),do(Z))=P(D|do(B),Z),$
if $(D\perp\!\!\!\perp Z|B)$ in $\mathcal{A}^{Z}_{B}$.
2. Insertion/Deletion of Actions:
$P(D|do(B),do(Z))=P(D|do(B)),$
if $(D\perp\!\!\!\perp Z|B)$ in $\mathcal{A}^{BZ}$.
We consider a causal graph with variables $Z$, $X$, and $\hat{Y}$, as shown in Figure 2. Starting with the law of total probability:
[…]
Step 1: Action/Observation Exchange. Using $(\hat{Y}\perp\!\!\!\perp X|Z)$ in $\mathcal{A}_{X}$, we apply the exchange rule:
$P(\hat{Y}|do(X),Z=z)=P(\hat{Y}|X,Z=z).$ (7)
Step 2: Insertion/Deletion of Actions. Using $(Z\perp\!\!\!\perp X)$ in $\mathcal{A}_{X}$, we simplify:
$P(Z=z|do(X))=P(Z=z).$ (8)
Substituting these into (6), we obtain:
[…]
The result can be compactly written as:
$p_{\Theta}(\hat{Y}|do(X))=\mathbb{E}_{p_{0}(Z)}\left[p_{\theta}(\hat{Y}|X,Z)\right],$ (10)
where $p_{0}(Z)$ denotes the prior distribution of environments. Do-calculus rules simplify interventional distributions by leveraging independence properties, enabling tractable objectives for causal inference.
A.3.2 Derivation for Variational Lower Bound
Below we show the derivation for the variational lower bound in Eq. (3).
$\quad\log\sum_{z}p_{\Theta}(\hat{Y}|X,Z=z)P(Z=z)$ (11)
$=\log\sum_{z}p_{\Theta}(\hat{Y}|X,Z=z)p(Z=z)\frac{q_{\phi}(Z=z|X)}{q_{\phi}(Z=z|X)}$
$\geq\sum_{z}q^{t+1}_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)\frac{p(Z=z)}{q_{\phi}(Z=z|X)}\quad$
(Jensen’s Inequality)
$=\sum_{z}q_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)-\sum_{z}q_{\phi}(Z=z|X)\log\frac{q_{\phi}(Z=z|X)}{p(Z=z)}$
$=\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]-KL\left(q_{\phi}(Z|X)\parallel p(Z)\right)$
(a) Previous Method
(b) Ours
Figure 8: Comparison between previous approach and our implementation.
A.3.3 Proof of Theorem 3.1
In this section, we provide the theoretical analysis of our optimization method, and an over is shown in Figure 8. Before proving Theorem 3.1, we prove the following theorem:
Theorem A.1.
Given the environment estimator before update $q_{\phi_{t}}$ and after update $q_{\phi_{t+1}}$, assuming the training process converges, then minimizing the KL divergence from $q_{\phi_{t+1}}$ to $q_{\phi_{t}}$, i.e., $\text{KL}(q_{\phi_{t+1}}||q_{\phi_{t}})$, is equivalent to applying an L2 norm on the parameters of the environment estimator $q_{\phi}$.

### v4

@@ 4.1 Experiment Setup
Datasets. For pre-training our model and PatchTST, we manually collected 17 distinct weekly-sampled diseases from Project Tycho (van Panhuis et al., 2018). For evaluation, we utilize five downstream datasets covering various diseases and locations: ILI USA (Centers for Disease Control and Prevention, 2023a), ILI Japan (National Institute of Infectious Diseases, 2023), COVID-19 USA (Dong et al., 2020), Measles England (Lau et al., 2020), and Dengue across countries (OpenDengue, 2023). Additionally, RSV (Centers for Disease Control and Prevention, 2023c) and MPox (Centers for Disease Control and Prevention, 2023b) infections in the US are used to test zero-shot performance. More details can be found in Appendix A.5.
Preprocessing. For pre-training datasets, we aggregate the time series based on time and locations to acquire the national-level infection trajectory. For all the datasets used in this study, we examine the infection trajectory for all diseases and locations, and filter the time series with extremely short observations or a large number of missing values to form a high-quality evaluation testbed. Then, we split the datasets into train/val/test sets and perform normalization on the time series.
Baselines We adopt baseline models from the comprehensive EpiLearn toolkit (Liu et al., 2024a), comparing our model against two categories: non-pretrained and pre-trained models. Non-pretrained baselines including RNN-based approaches like LSTM and GRU (Wang et al., 2020; Natarajan et al., 2023), the MLP-based model DLinear (Zeng et al., 2023), and transformer-based architectures (Wu et al., 2021; Zhou et al., 2021; 2022). Pre-trained baselines include state-of-the-art models such as PatchTST (Nie et al., 2022) and the time series foundation model MOMENT (Goswami et al., 2024). We provide further comparisons with PEM (Kamarthi & Prakash, 2023) in the few-shot setting and ARIMA (Panagopoulos et al., 2021) in the online forecasting setting in Appendix A.8. We perform hyperparameter tuning on the learning rate and weight decay for all models. ( See  A.6 for details)
Research Questions. In the following experiment, we propose and answer the following questions: Q1: How does the latent compartment contribute to epidemic forecasting? Q2: How does the model perform on diverse downstream datasets compared to other general time series models? Q3: How does the model perform in the few-shot setting with fewer observations? Q4: Does the proposed epidemic-aware regularization help? Q5: How does pre-training influence downstream performance?
4.2 Simulation on mechanistic model
Before conducting the empirical experiment, we first validate that CAPE’s latent compartments behave as intended under controlled simulations, which answers Q1. We constructed a simulation dataset based on the Susceptible-Infectious-Recovered-Deceased (SIRD) model for analysis. In this scenario, S, R, and D represent compartments with monotonic increasing or decreasing behaviors, while I shows a single peak pattern. We initialize our model with 16 latent compartments and assign S, I, R each with three compartment prototypes, D with two prototypes, and the rest without constraints. As shown in Figure 3, we observe that the trend and magnitude of latent compartment contribution roughly align with the actual compartments, which verifies the usefulness of our model.
Figure 2: Simulation on SIRD model. Left: Ground-truth trajectory; Right: Inferred compartment contributions.
Figure 3: Downstream performance vs. compartments distribution(See A.11).
@@ 4.3 Fine-Tuning (Full-Shot Setting)
To answer Q2, we finetune our model on diverse downstream datasets and compare performance across baselines from various designs. For non-pre-trained models, we train the entire model on the training split, while for pre-trained models, we fine-tune a few epochs on downstream datasets by transferring the task-specific head $h_{\psi}$ from pre-training to the forecasting task. We evaluate short-term and long-term performance by reporting mean MSE and MAE across horizons from 1 to 16 under 25 runs on test evaluation. From Table 1, we observe that CAPE achieves the best average performance across all downstream datasets compared with baselines.
4.4 Data Scarce Scenario, and Ablation Study
To answer Q3 and Q4, we evaluate the models in the few-shot and zero-shot regime with limited or no fine-tuning data, and perform ablations isolating the effects of pre-training and the epidemic-aware regularization.
Few-Shot Forecasting. Predicting outbreaks of new diseases or in unfamiliar locations is difficult for purely data-driven models with limited data, making few-shot and zero-shot forecasting essential. To simulate this, we reduce training data to [20%, 40%, 60%, 80%] and report average MSE over 1–16 time steps (Table 2). Key observations: (a) More training data consistently improves performance. (b) CAPE achieves the best results in most cases, showing strong few-shot capability. (c) Dlinear underperforms at 20% data compared to epidemic-pretrained models, but surpasses MOMENT on ILI USA and Measles when both are trained on 20%, highlighting the value of pre-training on epidemic time series. (More details are shown in  A.10).
@@ 4.4 Data Scarce Scenario, and Ablation Study
Dataset $\Delta$ (%) CAPE PatchTST MOMENT
ILI USA 9.26% 0.147 0.164 0.549
ILI Japan 17.06% 0.705 0.907 2.062
Measles 3.97% 0.145 0.167 0.533
MPox 20.00% 0.0004 0.0005 0.0013
Dengue (mixed) 10.17% 0.371 0.427 1.624
RSV 26.06% 0.834 1.128 1.849
Covid (daily interval) 13.80% 5.173 6.001 18.881
Table 4: Ablation study on removing pre-training and epidemic-aware regularization. Results are averaged over 25 runs of evaluation.
Dataset H w/o Pre-train w/o Epidemic Reg. CAPE
MSE MAE MSE MAE MSE MAE
ILI USA 1 0.148 0.114 0.180 0.144 0.174 0.139
2 0.229 0.151 0.200 0.145 0.192 0.141
4 0.409 0.202 0.297 0.171 0.299 0.171
8 0.575 0.241 0.565 0.240 0.469 0.221
16 0.640 0.289 0.652 0.278 0.650 0.278
Avg 0.400 0.199 0.379 0.196 0.357 0.190
ILI Japan 1 0.371 0.406 0.334 1.130 0.328 0.982
2 0.677 1.141 0.703 1.235 0.709 1.206
4 1.300 1.540 1.284 1.537 1.191 2.029
8 1.835 1.190 1.798 1.082 1.792 1.088
16 1.920 1.232 1.866 1.149 1.878 1.163
Avg 1.221 1.102 1.197 1.227 1.179 1.294
Zero-Shot Forecasting. We evaluate CAPE in a zero-shot setting by freezing transformer-based models with their pre-training heads. All models receive a 12-step historical input and predict the next 4 steps (Table 4). Key observations: (a) CAPE consistently outperforms baselines, confirming superior zero-shot ability. (b) Epidemic-specific pre-training yields better results than general pre-training (e.g., MOMENT), underscoring the importance of domain-specific data.
Ablation Study. We evaluate the contributions of CAPE’s components in Table 4. Removing pre-training leads to the largest degradation, with average MSE increasing from 0.357 to 0.400 on ILI USA. Also, dropping epidemic regularization hurts performance across most horizons and the average MSE (e.g., 1.179 to 1.197 on ILI Japan). Overall, CAPE consistently achieves the lowest errors, showing that both pre-training and epidemic regularization are important, with pre-training providing the greater benefit.
Figure 5: DBI between the embeddings of each pair of downstream datasets from the pre-trained model. (See Figure  8 for visualization.).
@@ 4.5 Deeper Analysis
To answer Q5, we (i) examine the representation quality, transferability, and robustness against distribution shift, and (ii) further explore the power of pre-training from two perspectives: compute budget and pre-training materials.
Transferability.
(a) Cross-Location: Pre-training on U.S. influenza data enables strong few-shot and zero-shot performance on Japanese influenza, highlighting the role of pre-training for regional generalization. (b) Cross-Disease: Despite excluding COVID-19 from pre-training, our model adapts well during downstream evaluation, outperforming the non–COVID-pretrained variant (Table 4) and surpassing MOMENT by 72.6%.
Representation Learning. To evaluate the representation quality learned during pre-training, we compute the DBI score of sample representations from different diseases. As shown in Figure 5 and  8, compared to PatchTST, CAPE is able to distinguish the representations from different diseases, effectively capturing the diverse underlying dynamics of different pathogens.
Tackling Distribution Shift. We define distribution shifts as changes in infection patterns from training to test data. We compute the Central Moment Discrepancy (CMD)(Zellinger et al., 2017) between training and test distributions for each disease. As shown in Figure 8, compared to the version without compartmental prototypes, CAPE achieves the lowest CMD scores, highlighting its effectiveness in mitigating distribution shifts.
Analysis on Pre-training. (a) Impact of Compute During Pre-training: Across four downstream datasets (Figure 3), more pre-training epochs consistently improve performance on Measles, with larger prototype sets $K$ yielding further gains. (b) Impact of Pre-Training Materials: Splitting pre-training data into respiratory vs. non-respiratory diseases (Figure 5) shows that alignment between pre-training and test disease types helps, though dataset size has a stronger effect. (c) Impact of Pre-Training Material Scale: Scaling the pre-training corpus (Figure 8) reveals a sharp performance jump at  60% reduction for both Measles and Dengue.
Figure 6: Downstream performance across pre-training ratios. More datasets are in  A.12.
Figure 7: Representation learned by CAPE vs. PatchTST after pretraining.
Figure 8: CMD scores w/ and w/o compartment estimate between train/test sets. (A.13)
@@ 5 Conclusion
We present CAPE, the first open-source pre-training framework for epidemic forecasting that learns flexible latent population states, termed compartmental prototypes, to address structural heterogeneity, hidden population states, and distribution shifts in epidemic pre-training. By designing large-scale epidemic self-supervised objectives with lightweight epidemic-aware regularization, CAPE captures transferable dynamics across diseases and regions. Extensive experiments demonstrate state-of-the-art generalization in zero-shot, few-shot, and cross-disease settings. In the future, we plan to extend CAPE to incorporate spatial dynamics for richer generalization, integrate principled approaches to uncertainty quantification, and further enhance the interpretability of the pre-trained model, ultimately advancing toward trustworthy and actionable epidemic forecasting.
Ethics Statement
We have adhered to the ICLR Code of Ethics in preparing this submission. This work does not involve human subjects, personally identifiable data, or sensitive information. All datasets used are publicly available benchmark datasets, and we follow their respective usage and licensing guidelines. The proposed methods are designed for advancing research in high-dimensional time series forecasting and do not raise foreseeable risks of harm.
Reproducibility Statement
We provide an anonymous repository containing the full source code and implementation details of our proposed CAPE at https://anonymous.4open.science/r/CAPE_ICLR26-A041/. Detailed descriptions of model architectures, training protocols, and hyperparameters are included in the main text and appendix. These resources are intended to ensure that all reported results can be independently reproduced.
@@ A.1 Potential Impact
Early Insights. We provide novel insights into how pre-training and compartment modeling improve epidemic forecasting. Our results demonstrate that pre-training significantly enhances model accuracy, with gains increasing as more pre-training data is incorporated. This finding paves the way for future research to develop foundational models in epidemic analysis using larger datasets. Additionally, we confirm the importance of accounting for both inherent disease dynamics and compartmental factors to achieve robust forecasting performance.
@@ A.2 Limitations
While CAPE shows strong performance, it has several limitations. First, it does not currently incorporate uncertainty estimation, which is important for risk-aware decision-making. Second, its effectiveness may be constrained by the scale and diversity of available data. Lastly, the learned compartment representations lack interpretability, limiting their transparency and potential for insight in public health contexts.
A.3 Use of LLMs
We use LLMs to check grammar and polish the language of this paper for clarity.
A.4 Theoretical Analysis
A.4.1 Inference of $R_{0}$
(1) Disease-Free Equilibrium (DFE): Calculation of $\mathbf{F}$ and $\mathbf{V}$ requires defining the population flow, and since our framework does not utilize explicit population flows, we use the learned latent compartment contributions, $\bm{\pi}$, as a potential differentiable proxy that correlates with population, as demonstrated in Figure 3. Specifically, we establish the compartment contributions at DFE: $\bm{\pi}^{*}$, which represents a scenario with no ongoing epidemic, by feeding a zero-infection time series, $\mathbf{X}_{\text{DFE}}\in\mathbb{R}^{T\times 1}$, into the encoder:
$\bm{\pi}^{*}=\frac{1}{C}\sum_{c=1}^{C}\text{softmax}(\bm{E}_{\text{DFE}}\mathbf{E}^{T}[c,:])\in\mathbb{R}^{K},\\ \quad\bm{E}_{\text{DFE}}\leftarrow f_{\text{enc}}(\mathbf{X}_{\text{DFE}}),$ (7)
where $\bm{E}_{\text{DFE}}$ are the weighted sum of latent compartment prototypes at the DFE.
(2) Calculation of $\mathbf{F}$: The $\mathbf{F}$ matrix, which quantifies new infections, is estimated by applying perturbation on each infectious compartment: $\mathbf{H}_{\text{pert},(j)}=\phi(\hat{\bm{\pi}},\mathbf{E}_{\text{DFE}})=(1-\alpha)\mathbf{E}_{\text{DFE}}+\alpha\hat{\bm{\pi}}^{T}\mathbf{E}$, where $\alpha=0.1$, $\hat{\bm{\pi}}=\frac{\bm{\pi}^{*}+\bm{\epsilon}_{j}}{\|\bm{\pi}^{*}+\bm{\epsilon}_{j}\|_{1}}$ and $\bm{\epsilon}_{j}$ is a perturbation applied on the entry (compartment) $j$. Therefore, we get the element of $\mathbf{F}$ via:
$F_{ij}=\max(0,\pi^{\text{new},(j)}_{i}-\pi^{*}_{i}),\\ \quad\bm{\pi}^{\text{new},(j)}=g(\mathbf{H}_{\text{pert},(j)}).$ (8)
(3) Calculation of $\mathbf{V}$: $\mathbf{V}$ characterizes transition rates out of infectious compartments, where $V_{jj}$ represents the total departure rate from compartment $j$ and $V_{ij}$ captures transitions from compartment $j$ to compartment $i$. For each infectious compartment $j$, we initialize with unit mass $\bm{\pi}^{\text{unit, (j)}}$ and acquire the updated contributions in a similar way: $\bm{\pi}^{\text{evolved},(j)}=g(\phi(\bm{\pi}^{\text{unit, (j)}},\mathbf{E}_{\text{DFE}}))$. Then the V matrix elements are computed as:
$V_{ij}=\begin{cases}\max(0,\pi^{\text{evolved},(j)}_{i})&\text{if }i\neq j,\\ \pi_{j}^{\text{unit},(j)}-\pi^{\text{evolved},(j)}_{j}&\text{if }i=j,\end{cases}$ (9)
where $V_{ij}$ measures mass appearing in compartment $i$ when $j$ initially contains unit mass, and $V_{jj}$ quantifies the total rate of departure from compartment $j$. Finally, a linear layer is applied to align the scale of the estimated lower- and upper-bound with the ground-truth range of $R_{0}$ and the loss is computed as: $\mathcal{L}_{R_{0}}=\text{RELU}(\hat{R}^{\text{lower}}_{0},R^{\text{lower}}_{0})+\text{RELU}(\hat{R}^{\text{upper}}_{0},R^{\text{upper}}_{0})$. This NGM-based approach provides a theoretically grounded method for computing $R_{0}$ that respects the compartmental structure while being differentiable for end-to-end training.
A.4.2 Proof of Equation  5
Setup and Preliminaries. Let $V\in\mathbb{C}^{m\times n}$. The (operator) $2$-norm of a matrix $M$ is
$\left\lVert M\right\rVert_{2}\;=\;\sup_{x\neq 0}\frac{\left\lVert Mx\right\rVert_{2}}{\left\lVert x\right\rVert_{2}}\;=\;\max_{\left\lVert x\right\rVert_{2}=1}\left\lVert Mx\right\rVert_{2}.$
The singular values $\sigma_{1}(V)\geq\cdots\geq\sigma_{r}(V)\geq 0$ (with $r=\operatorname{rank}(V)$) are, by definition, the nonnegative square roots of the eigenvalues of $V^{\ast}V$ (where denotes conjugate transpose), counted with multiplicity and ordered nonincreasingly.
For any matrix $M$ and any vector $x$,
$\sigma_{\min}(M)\,\|x\|_{2}\;\leq\;\|Mx\|_{2}\;\leq\;\sigma_{\max}(M)\,\|x\|_{2},\quad\|M\|_{2}=\sigma_{\max}(M).$
For square $M$, the singular values $\{\sigma_{i}(M)\}_{i=1}^{n}$ and eigenvalues $\{\lambda_{i}(M)\}_{i=1}^{n}$ satisfy
$\prod_{i=1}^{n}\sigma_{i}(M)=\sqrt{\det(M^{*}M)}=|\det M|=\prod_{i=1}^{n}|\lambda_{i}(M)|.$
In particular, by comparing geometric means to extrema,
$\sigma_{\min}(M)\;\leq\;\Big(\prod_{i=1}^{n}\sigma_{i}(M)\Big)^{1/n}=\Big(\prod_{i=1}^{n}|\lambda_{i}(M)|\Big)^{1/n}\;\leq\;\max_{i}|\lambda_{i}(M)|=\rho(M).$
Thus, for any square $M$,
$\sigma_{\min}(M)\;\leq\;\rho(M)\;\leq\;\sigma_{\max}(M)=\|M\|_{2}.$
Theorem A.1 (Bounds for $R_{0}$).
Let $F,V\in\mathbb{C}^{n\times n}$ with $V$ invertible, and define
$R_{0}\;\equiv\;\rho(FV^{-1}),$
where $\rho(\cdot)$ denotes the spectral radius, $\|\cdot\|_{2}$ the operator $2$-norm, and $\sigma_{\max}(\cdot),\sigma_{\min}(\cdot)$ the maximal and minimal singular values, respectively. Then $R_{0}$ satisfies the bounds
$\boxed{\;\frac{\sigma_{\min}(F)}{\sigma_{\max}(V)}\;\leq\;\rho(FV^{-1})\;\leq\;\frac{\sigma_{\max}(F)}{\sigma_{\min}(V)}.\;}$
Derivation for Lower bound.